Controls Meeting 26-Jan-2012

From GlueXWiki
Jump to: navigation, search


  • Announcements
  • Minutes from previous meeting 15-Dec-2011
  • Counting house status - Hovanes
    • UPS, networking, rack room, counting room, office area, phones, printers, furniture, etc.
  • Equipment orders - Mark
  • Solenoid_Controls_Redesign - Mark, Josh, Scot, Yi, Elliott
    • QD, shunt, ground fault detector, racks, UPS power to racks
    • coil inductances, schematics, etc. - Scot
    • fast/slow DAQ integration, Labview licenses - Yi, Elliott
  • Experiment Control Supervisor - Vardan, Hovanes, Elliott
  • EPICS - Hovanes, Nerses
    • archiver, alarm system, scalers
  • Test stand - Hovanes, Josh
  • Target, goniometer, other new systems - Hovanes, Elliott
  • PLC programming - Josh
    • A-B class, program redesign, new features, new computer
  • Elogs and databases - Elliott
    • Midas elog
    • Operator elog
    • inventory database - IRMIS, CLAS, other?
  • Mantis task tracker


9:30 AM Thurs 26-Jan-2012 CC L207


Next Meeting

9:30am 9-Feb-2012 CC L207


Present: Elliott W, Beni Z, Josh B, Mark S, Hovanes E, Scot S.

Counting house status

Hovanes described recent plans and progress concerning outfitting the counting house. Interim networks and phones are in and the UPS installation is being planned along with routing of clean and UPS power. Next in is racks and our network equipment. He is also looking into office area furniture, displays and remote conferencing equipment. At some point we'll have to design and outfit the control room.

Equipment orders

  • Mark has ordered some tools, test equipment and rack hardware, but has more to go.
  • Hovanes ordered a 5 TB Dell RAID system.
  • Hovanes ordered a 8-axis Newport motor controller, likely to be used on a number of beamline components.
  • Three VME cpu's arrived, Beni will use them, 3 more on the way.

Solenoid controls

Elliott reported that we are still on schedule for magnet cooldown in Nov-2012. There is a great dealt to do so we cannot waste time. A working control system needs to be designed, purchased, installed and tested by this summer.

  • There are no known showstoppers, but we still need the results of the FMEA from George to know the level of redundancy needed.
  • The QD is under control.
  • It now appears that Yi's National Instruments PXI solution solves the fast/slow daq and isolation problems.
  • Labview licenses don't look to be particularly expensive.
  • Scot noted we may have more coil tap signals than expected since George wants taps in more places.
  • Scot ordered another Lakeshore temperature monitor. We have no spare channels in the four units.


  • Nerses moving from scaler readout to HV/LV control, including CAEN, Wiener and CANBus systems.
  • Hovanes waiting to hear from Accelerator on archivers and integration with CSS.
  • Hovanes started working on an alarm system (BEAST from SNS).


Josh just returned from a very informative advanced class on A-B PLC programming. He learned a lot about program organization, programming techniques, languages, connections, etc. Tim recognizes that the program redesign effort will take a lot of Josh's time over the next few months. Due to network security constraints we'll need to place a new computer in the Counting House to host the A-B software.

Elogs and databases

Elliott reported we will meet in two weeks with Accelerator concerning logbooks and other items. The Midas elog system is working well but everyone needs to start entering manuals, procedures and use it for maintenance log. Josh suggested we create a materials certification logbook, Elliott will do this if Josh sends him the required columns. Everyone recognizes we need a components database but it is not clear which to use. IRMIS will be released in Feb/Mar. We may have to design our own if IRMIS doesn't appear or doesn't fit.